# 6G: The Intelligent Network of Everything – A Comprehensive Vision, Survey, and Tutorial

### v2

@@ V-A Communication-Oriented Use Cases
  • Extreme Time-Sensitive and Mission-Critical Communications: This use case is intended to support time-sensitive and mission-critical applications in versatile environments, such as smart factories, smart hospitals, autonomous vehicle systems, and emergency operations. Extreme reliability and very low latency are required since there is no room for severe errors. The main enablers for extreme reliability include network- and link-level diversity methods, high channel coding rates, robust modulation orders, and advanced retransmission mechanisms. Extremely low latency can be attained through flat higher-layer management procedures, bringing content close to end-devices, minimized control signaling, fast network access mechanisms, and short transmission frames. Since reliability and low latency are conflicting requirements, there is always a trade-off between them. Note that this use case can be seen as an extension of 5G URLLC.
@@
  • Ultra-Massive Communications: This use case aims to support an ultra-massive number of different types of devices in diverse application scenarios. The range of devices is wide, including IoT sensors, devices, wearables, machines, robots, vehicles, drones, etc. Application scenarios vary from smart environments (e.g., cities, factories, hospitals, buildings, warehouses, etc.) to wide-area IoT (e.g., environmental/industrial monitoring) and low-power sensor networks. Due to versatile devices and applications, many performance dimensions need to be considered, such as connection density, coverage, mobility, capacity, and reliability. The key enablers for massive connectivity include flexible spectrum operation, fine-grained frame structure, massive grant-free network access, robust transmission techniques, and efficient resource management. This use case is an extension of 5G mMTC.
@@
  • Global-Scale Communications: The goal of the use case is to provide worldwide super-coverage, including remote areas on the land, at the sea, and in the air. This use case supports diverse global-level application areas, including remote mobile internet, global IoT, environmental monitoring, industrial tracking, and maritime/aerial communications, to mention a few. Aligned with the United Nations Sustainable Development Goals, digital divide may be alleviated by providing decent-quality communications to remote and developing areas around the world. Global coverage is enabled by integrated terrestrial and non-terrestrial networks that exploit satellite, aerial, and mobile communications.
@@
  • Hyper-Mobility Communications: This use case is meant to support communications at high velocities, even up to 500–1000 km/h. Example application scenarios are high-velocity transportation systems, such as high-speed trains. The main enablers include flexible operation within a suitable spectrum range, fast handovers, and enhanced mobility management techniques. In general, lower frequencies are better for higher-mobility scenarios due to their wider coverage and less frequent handovers.
TABLE II: Main performance requirements for 6G
@@
V-B Beyond-Communication-Oriented Use Cases
@@
  • Network Intelligence: The target of this use case is to provide AI/ML-assisted intelligent network services and applications, such as intelligence as a service, network-assisted intelligent systems (industry, transportation, health, etc), computation offloading, and edge caching. This use case requires powerful computation and storage abilities, advanced AI/ML methods, abundant data acquisition, and efficient AI/ML model training/inference.
  • Network Sensing: This use case aims to expand the capabilities of cellular networks to sensing and positioning. Example applications include object identification, shape recognition, activity detection, range estimation, velocity evaluation, and movement tracking. This use case is enabled by the integration of communication and sensing.
  • Network Energy: The aim of this use case is to support wireless energy transfer (WET) services to power network nodes, such as IoT sensors and low-power devices. In particular, wireless energy services aim to advance the long life-time, low-complexity, autonomy, mobility, and novel usages of low-energy networks. This use case can provide customized energy services to diverse vertical industries. Energy beamforming is a key technology to provide wireless energy capabilities.
VI Main Performance Requirements for 6G
In this section, the main performance requirements are discussed for 6G. In November 2023, ITU-R introduced 15 performance capabilities for IMT-2030 in Recommendation ITU-R M.2160-0, i.e., ”IMT-2030 Framework” [36]. The numerical target values were defined for nine of them. In the following, we review these target requirements and compare them to those of IMT-2020 described in Recommendation ITU-R M.2083-0 [11]. In addition, we present some typical values proposed in the 6G literature (see [88], and the references therein). The target requirements are summarized in Table II. At the end of this section, we briefly list the rest of the IMT-2030 capabilities identified by ITU-R. Note that ITU-R has not defined any target values for these capabilities since most of them are difficult to measure numerically.
@@ VI Main Performance Requirements for 6G
  • Area Traffic Capacity: The area traffic capacity is defined as the total aggregated throughput per unit area. Compared to IMT-2020’s 10 Mbit/s/m2, ITU-R recommends the values of 30 and 50 Mbit/s/m2 for IMT-2030, while greater values could also be examined. The values typically presented in the 6G literature are much higher than those of ITU-R, i.e., 1 and 10 Gbit/s/m2.
  • Latency: The user plane latency is the time spent by the network from sending a packet to receiving it, assuming unloaded conditions. IMT-2020’s latency requirement is 1 ms. According to ITU-R, the target values could range from 0.1 to 1 ms for IMT-2030. The target of 0.1 ms is typical in the 6G literature.
@@
VII AI/ML for 6G
AI/ML is expected to play a revolutionary role in 6G networks. The extensive use of AI/ML can fundamentally change the way how mobile networks are designed, operated, and managed. In this section, we provide a brief introduction to the fundamentals of AI/ML and review three promising ML methods for 6G.
VII-A Introduction to AI and ML
@@ VII-A1 Artificial Intelligence
The potential of AI was recognized early on, but with unrealistic optimism in practical implementations. After decades of great expectations, the use of AI finally boomed in the 2010s. This was due to advances in key areas, such as computer technology, ML, and big data. Since the 2010s, the world’s top technology companies have invested heavily in AI. A detailed history of AI is presented in [172]. Today, AI is used everywhere. Through a vast variety of applications, AI has penetrated deeper into the different levels of modern society. Typical applications include Internet search engines, recommendation systems, online advertising, social media, entertainment, gaming, voice assistants, smart phones, smart home devices/appliances, online food delivery, electric scooter services, assisted/self-driving cars, drones, robots, etc. In the future, the use of AI will significantly increase in all areas of life and society.
VII-A2 Machine Learning
@@ VII-A3 Main Types of ML
TABLE III: Summary of ML methods for 6G
@@
VII-A4 Main Types of Training
In ML, there are two main types of training: offline and online [177].
Figure 10: Key ML methods for 6G.
VII-B Key ML Methods for 6G
In this section, we discuss three ML methods that can provide diverse benefits to 6G networks. These methods include DL, federated learning (FL), and transfer learning (TL), as illustrated in Figure 10. The opportunities and challenges of these methods are summarized in Table III.
VII-B1 Deep Learning
@@ VII-B3 Transfer Learning
TABLE IV: Summary of spectrum-level technologies for 6G
@@ VIII Potential Technologies for 6G
VIII-A Spectrum-Level Technologies for 6G
To fulfill the demanding performance, service, and application requirements of 6G, it is vital to support a massive frequency range from sub-6 GHz and centimeter wave (cmWave) to mmWave and THz bands. At the cost of smaller system bandwidths, lower carrier frequencies with favorable propagation characteristics are preferable for wider coverage and higher mobility scenarios, such as rural areas, wide-area IoT, and high-mobility systems. At the cost of harsher propagation conditions, higher frequencies with larger bandwidths enable higher capacity and throughput in denser network scenarios, such as hotspot, urban, and sub-urban areas. In this picture, THz communication is seen as a promising technology to achieve extremely high data rates. In addition, optical wireless communication (OWC) is a potential candidate to offer complementary frequency resources for special scenarios, such as short-range indoor access and long-range outdoor backhaul links. These two technologies are summarized in Table IV and discussed in detail below.
@@ VIII-A1 THz Communications
  • Past and Present: In the literature, THz research on mobile networks started to gain increasing interest around the mid-2010s [211]. However, back then, the main focus of higher frequency research was on mmWave communication for 5G. Since 5G, with mmWave operation as a key technology, was launched in the late 2010s, the worldwide research focus was shifted to THz communications for 6G. Currently, academia is extensively studying THz communication from a wide range of perspectives, from theoretical aspects and channel modeling to transceiver designs, assisting technologies, and performance evaluations [209]. In addition, cooperation between academia and industry is narrowing the gap between theory and practice, producing fruitful ideas and leading to novel designs and demonstrations.
@@
    In the spectrum regulatory domain, WRC 2019 defined the conditions for the use of a spectrum ranging from 275 to 450 GHz by land mobile and fixed services [212]. As a result, there is currently a spectrum of 160 GHz for THz communications, with no specific conditions to protect Earth exploration satellite services. Further spectrum policies for THz communications will be defined in the future WRCs. In the standardization domain, IEEE specified the first standard (IEEE Std. 802.15.3d-2017) for fixed wireless point-to-point THz communications at 252-321 GHz frequencies, with target applications of wireless backhaul/fronthaul, wireless links in data centers, kiosk downloading, and intra-device communications [213].
    Although there are no standards for mobile THz communications, IEEE Std. 802.15.3d-2017 will pave the way toward 6G THz standardization. In 2021, the IEEE Communications Society’s Radio Communications Committee established a special interest group on THz communications to advance research, development, and standardization activities toward 6G and beyond [209]. In the existing mobile networks, 5G is the first generation that operates at frequencies higher than 6 GHz. Specifically, 5G supports mmWave communications at two frequency ranges, i.e., FR2-1: 24.5-52.6 GHz (Release 15 [1]) and FR2-2: 52.7-71 GHz (Release 17 [17]). 6G will continue this trend by going beyond 100 GHz, becoming the first generation to enter THz frequencies.
@@
    There are many challenges in integrating THz communication as a key part of 6G networks and freeing its great potential. The main challenges include high propagation losses (spreading and molecular absorption losses), easy blockages, severe hardware impairments (oscillator phase noise and RF non-linearity), and complex hardware design (signal generation and detection) [214]. Due to the high molecular absorption loss (especially above 800 GHz), there are large chunks of spectrum in the THz band that may be unsuitable for mobile communications [214]. Other parts of the spectrum, where attenuation is tolerable, are called transmission windows and considered more preferable for 6G purposes [214]. There are massive amounts of spectrum in these windows, ranging from tens to even hundreds of GHz.
@@
    A major challenge is how to handle hardware impairments, such as phase noise and power amplifier non-linearity, as they become more severe at higher frequencies. To a certain extent, a wider SCS can alleviate hardware impairments in multi-carrier-based systems. Another solution is to employ single-carrier waveforms for the operation at very high frequencies since they are more robust against phase noise and other hardware impairments [216]. In terms of transceiver hardware, signal generation and detection are particularly challenging for THz communications [210], and novel solutions are required for cost-, power-, and energy-efficient communications. As THz communication is prone to blockages, a possible solution is to utilize RIS technology, which can reflect the transmitted signals to bypass blockages [217, 218]. With controlled reflections, RISs can also be used to alleviate the distance problem by extending the coverage of THz communications [217, 218].
  • Literature and Future Directions: In the literature, THz mobile communication has been actively studied since the mid-2010s, and it has been one of the major 6G topics since the late 2010s. The THz literature covers a broad range of topics, such as device technologies, channel modeling, communication designs, networking aspects, application scenarios, regulation/standardization, and experiments [209]. The latest THz research has been comprehensively reviewed in numerous studies [219, 214, 220, 221, 222, 210, 213, 223, 218, 224, 225, 226, 217, 227, 228, 209, 229, 230, 231, 232, 233, 234, 235, 236, 237, 238, 239, 240, 241].
    In [219], a survey was conducted on THz communications, reviewing key transceiver technologies, system designs, applications, and challenges. The work in [214] provided a comprehensive study on wireless communications above 100 GHz, with the main focus on THz applications, transceiver architectures, beamforming designs, and channel properties toward 6G and beyond. In [220], a short survey discussed the significance, standardization, scenarios, applications, and open problems of THz communications in the context of beyond 5G networks. THz communication systems were reviewed in [221] from the perspective of signal generation, channel modeling, contender technologies, applications, standardization, and future directions.
    In [222], the THz band was considered for sensing, imaging, and localization applications, especially in the context of 6G use cases. THz ultra-massive MIMO was comprehensively examined in [210] and [227]. Standardization for sub-THz communications in IEEE 802.15.3d was discussed in [213]. 6G THz networks were studied from the full-communication stack perspective in [223], addressing link- and system-level challenges. RIS-assisted THz communication was explored for 6G networks in [218, 217, 235]. In [224], a tutorial was provided on signal processing techniques for THz communications, with an emphasis on ultra-massive MIMO and RIS systems.
@@
VIII-A2 Optical Wireless Communications

### v3

@@ VI-A COMMUNICATION-ORIENTED USE CASES
  • Extreme Time-Sensitive and Mission-Critical Communications: This use case is intended to support time-sensitive and mission-critical applications in versatile environments, such as smart factories, smart hospitals, autonomous vehicle systems, and emergency operations. Extreme reliability and ultra-low latency are required since there is no room for severe errors. The main enablers for extreme reliability include network- and link-level diversity methods, high channel coding rates, robust modulation orders, and advanced retransmission mechanisms. Extremely low latency can be attained through flat higher-layer management procedures, bringing content close to end-devices, minimized control signaling, fast network access mechanisms, and short transmission frames. Since reliability and low latency are conflicting requirements, there is always a trade-off between them. Note that this use case can be seen as an extension of 5G URLLC.
@@
  • Ultra-Massive Communications: This use case aims to support an ultra-massive number of different types of devices in diverse application scenarios. The range of devices is wide, including IoT sensors, devices, wearables, machines, robots, vehicles, drones, etc. Application scenarios vary from smart environments (e.g., cities, factories, hospitals, buildings, warehouses, etc.) to wide-area IoT (e.g., environmental/industrial monitoring) and low-power sensor networks. Due to versatile devices and applications, many performance dimensions need to be considered, such as connection density, coverage, mobility, capacity, and reliability. The key enablers for massive connectivity include flexible spectrum operation, fine-grained frame structure, massive grant-free network access, robust transmission techniques, and efficient resource management. The challenge is to build such a flexible network. This use case is an extension of 5G mMTC.
@@
  • Global-Scale Communications: The goal of the use case is to provide worldwide super-coverage, including remote areas on the land, at the sea, and in the air. This use case supports diverse global-level application areas, including remote internet, global IoT, environmental monitoring, industrial tracking, and maritime/aerial communications, to mention a few. Aligned with the United Nations Sustainable Development Goals, digital divide may be alleviated by providing internet access to remote and developing areas around the world. Moreover, we can fight climate change with worldwide environmental monitoring capabilities. Global coverage is enabled by integrated terrestrial and non-terrestrial networks that exploit satellite, aerial, and mobile communications.
@@
  • Hyper-Mobility Communications: This use case is meant to support communications at high velocities, even up to 500–1000 km/h. Example application scenarios are high-velocity transportation systems, such as ultra-high-speed trains. The main enablers include flexible operation within a suitable spectrum range, fast handovers, and enhanced mobility management techniques. In general, lower frequencies are better for higher-mobility scenarios due to their wider coverage and less frequent handovers.
TABLE IV: Main performance requirements for 6G
@@
VI-B BEYOND-COMMUNICATION-ORIENTED USE CASES
@@
  • Network Intelligence: The target of this use case is to provide AI/ML-assisted intelligent network services and applications, such as intelligence as a service, network-assisted intelligent systems (industry, transportation, healthcare, etc), computation offloading, and edge caching. This use case requires powerful computation and storage abilities, advanced AI/ML methods, abundant data acquisition, and efficient AI/ML model training/inference. A major challenge is to provide a ubiquitous communication, computation, and caching infrastructure.
  • Network Sensing: This use case aims to expand the capabilities of cellular networks to sensing and positioning. Example applications include object identification, shape recognition, activity detection, range estimation, velocity evaluation, and movement tracking. Typical application environments include industrial and healthcare. For example, network sensing can be used to detect a gas leakage in a factory or detect a person falling down in an elderly care with an automatic call for help. This use case is enabled by the integration of communication and sensing. The main challenge is to obtain a high accuracy in sensing.
  • Network Energy: The aim of this use case is to support wireless energy transfer (WET) services to power network nodes, such as IoT sensors and low-power devices. In particular, wireless energy services aim to advance the long life-time, low-complexity, autonomy, mobility, and novel usages of low-energy networks. This use case can provide customized energy services to diverse vertical industries. Energy beamforming is a key technology to provide wireless energy capabilities. In energy beamforming, typical challenges are to achieve high energy transfer efficiency and sufficiently long transfer ranges.
VII MAIN PERFORMANCE REQUIREMENTS FOR 6G
In this section, the main performance requirements are discussed for 6G. In November 2023, ITU-R introduced 15 performance capabilities for IMT-2030 in Recommendation ITU-R M.2160-0, i.e., ”IMT-2030 Framework” [20]. The numerical target values were defined for nine of them. In the following, we review these target requirements and compare them to those of IMT-2020 described in Recommendation ITU-R M.2083-0 [10]. In addition, we present some typical values proposed in the 6G literature (see [99], and the references therein). The target requirements are summarized in Table IV. At the end of this section, we briefly list the rest of the IMT-2030 capabilities identified by ITU-R. Note that ITU-R has not defined any target values for these capabilities since most of them are difficult to measure numerically.
@@ VII MAIN PERFORMANCE REQUIREMENTS FOR 6G
  • Area Traffic Capacity: The area traffic capacity is defined as the total aggregated throughput per unit area. Compared to 10 Mbit/s/m2 of IMT-2020, ITU-R recommends the values of 30 and 50 Mbit/s/m2 for IMT-2030, while greater values could also be examined. The values typically presented in the 6G literature are much higher than those of ITU-R, i.e., 1 and 10 Gbit/s/m2.
  • Latency: The user plane latency is the time spent by the network from sending a packet to receiving it, assuming unloaded conditions. The latency requirement of IMT-2020 is 1 ms. According to ITU-R, the target values could range from 0.1 to 1 ms for IMT-2030. The target of 0.1 ms is typical in the 6G literature.
@@
VIII AI/ML FOR 6G
In this section, we provide a brief introduction to the fundamentals of AI/ML, review three promising ML methods for 6G, and finally have a high-level discussion on the opportunities and challenges of AI/ML in 6G networks.
VIII-A INTRODUCTION TO AI AND ML
@@ VIII-A1 ARTIFICIAL INTELLIGENCE
The potential of AI was recognized early on, but with unrealistic optimism in practical implementations. After decades of great expectations, the use of AI finally boomed in the 2010s. This was due to advances in key areas, such as computer technology, ML, and big data. Since the 2010s, the top technology companies have invested heavily in AI. A detailed history of AI is presented in [190]. Today, AI is used everywhere. Through a vast variety of applications, AI has penetrated deeper into the different levels of modern society. Typical applications include Internet search engines, recommendation systems, online advertising, social media, entertainment, gaming, voice assistants, smart phones, smart home devices/appliances, online food delivery, electric scooter services, assisted/self-driving cars, drones, robots, etc. In the future, the use of AI will significantly increase in all areas of life and society.
VIII-A2 MACHINE LEARNING
@@ VIII-A3 MAIN TYPES OF ML
VIII-A4 MAIN TYPES OF TRAINING
In ML, there are two main types of training: offline and online [195].
Figure 10: Key ML methods for 6G.
TABLE V: Summary of ML methods for 6G
@@ VIII-A4 MAIN TYPES OF TRAINING
VIII-B KEY ML METHODS FOR 6G
In this section, we discuss three ML methods that can provide diverse benefits to 6G networks. These methods include DL, federated learning (FL), and transfer learning (TL), as illustrated in Figure 10. The opportunities and challenges of these methods are summarized in Table V.
VIII-B1 DEEP LEARNING
@@ VIII-B3 TRANSFER LEARNING
VIII-C DISCUSSION
AI/ML is expected to play a revolutionary role in 6G networks. The extensive use of AI/ML can fundamentally change the way how mobile networks are designed, operated, and managed. While AI/ML holds great potential when integrated into 6G networks, it also brings significant challenges across technological, societal, and ethical domains. To ensure an efficient integration of AI/ML into 6G networks, it requires overcoming substantial technological obstacles in developing a ubiquitous and powerful computing infrastructure from the core to the edge, advanced data management processes, and practical data acquisition methods. In particular, real-time processing in AI/ML-enhanced network environments is a major challenge. Since the extensive use of AI/ML requires substantial amounts of energy, developing sustainable and energy-efficient AI/ML solutions is crucial. These technological challenges are discussed in detail in Section IX-E.
As pervasive AI/ML requires vast amounts of data, possibly personal or sensitive, to be collected and analyzed, it raises concerns on the privacy, data ownership, and consent of the users. In data ownership, the question is who owns the data, is it the users who generate it, the companies who manage it, or the governments who regulate it. This is a critical question since the ownership affects how data is being used. Due to the lack of transparency in AI/ML systems, it may be difficult for users to fully understand how their data is being used. Thus, getting a genuine consent from the users to use their personal information becomes challenging. The ethical considerations of AI/ML in decision making revolve around transparency and accountability. If there is a lack of transparency, it can be challenging to pinpoint who is accountable when errors occur. Clear regulations are needed to protect the privacy of the users, prevent the misuse of the data, and ensure the responsible use of AI/ML in the 6G era. Further details on responsible AI for mobile networks can be found in [227].
@@ VIII-C DISCUSSION
TABLE VI: Summary of spectrum-level technologies for 6G
@@ IX POTENTIAL TECHNOLOGIES FOR 6G
IX-A SPECTRUM-LEVEL TECHNOLOGIES FOR 6G
To fulfill the demanding performance, service, and application requirements of 6G, it is vital to support a massive frequency range from sub-6 GHz and centimeter wave (cmWave) to mmWave and THz bands. At the cost of smaller system bandwidths, lower carrier frequencies with favorable propagation characteristics are preferable for wider coverage and higher mobility scenarios, such as rural areas, wide-area IoT, and high-mobility systems. At the cost of harsher propagation conditions, higher frequencies with larger bandwidths enable higher capacity and throughput in denser network scenarios, such as hotspot, urban, and sub-urban areas. In this picture, THz communication is seen as a promising technology to achieve extremely high data rates. In addition, optical wireless communication (OWC) is a potential candidate to offer complementary frequency resources for special scenarios, such as short-range indoor access and long-range outdoor backhaul links. These two technologies are summarized in Table VI and discussed in detail below.
@@ IX-A1 THZ COMMUNICATIONS
  • Past and Present: In the literature, THz research on mobile networks started to gain increasing interest around the mid-2010s [230]. However, back then, the main focus of higher frequency research was on mmWave communication for 5G. Since 5G, with mmWave operation as a key technology, was launched in the late 2010s, the worldwide research focus was shifted to THz communications for 6G. Currently, academia is extensively studying THz communication from a wide range of perspectives, from theoretical aspects and propagation characteristics to transceiver designs, assisting technologies, and performance evaluations [228]. In addition, cooperation between academia and industry is narrowing the gap between theory and practice, producing fruitful ideas and leading to novel designs and demonstrations.
@@
    In the spectrum regulatory domain, WRC 2019 defined the conditions for the use of a spectrum ranging from 275 to 450 GHz by land mobile and fixed services [231]. As a result, there is currently a spectrum of 160 GHz for THz communications, with no specific conditions to protect Earth exploration satellite services. As discussed in Section I-C2, at WRC 2023, ITU-R identified five frequency bands above 100 GHz to be further studied for IMT-2030, with potential approval at WRC 2031 [21]. These bands were 102–109.5 GHz, 151.5–164 GHz, 167–174.8 GHz, 209–226 GHz and 252–275 GHz. ITU-R has also examined the feasibility of mobile communications at THz frequencies, concluding that it is feasible to operate at the spectrum above 100 GHz [22]. Further spectrum policies for THz communications will be defined in the future WRCs. In the standardization domain, IEEE specified the first standard (IEEE Std. 802.15.3d-2017) for fixed wireless point-to-point THz communications at 252-321 GHz frequencies, with target applications of wireless backhaul/fronthaul, wireless links in data centers, kiosk downloading, and intra-device communications [232].
    Although there are no standards for mobile THz communications, IEEE Std. 802.15.3d-2017 will pave the way toward 6G THz standardization. In 2021, the IEEE Communications Society Radio Communications Committee established a special interest group on THz communications to advance research, development, and standardization activities toward 6G and beyond [228]. In the existing mobile networks, 5G is the first generation that operates at frequencies higher than 6 GHz. Specifically, 5G supports mmWave communications at two frequency ranges, i.e., FR2-1: 24.5-52.6 GHz (Release 15 [1]) and FR2-2: 52.7-71 GHz (Release 17 [15]). 6G will continue this trend by going beyond 100 GHz, becoming the first generation to enter THz frequencies.
@@
    There are many challenges in integrating THz communication as a key part of 6G networks and freeing its great potential. The main challenges include high propagation losses (spreading and molecular absorption losses), easy blockages, severe hardware impairments (oscillator phase noise and RF non-linearity), complex hardware design (signal generation and detection), and high power consumption of THz devices [233]. Due to the high molecular absorption loss (especially above 800 GHz), there are large chunks of spectrum in the THz band that may be unsuitable for mobile communications [233]. Other parts of the spectrum, where attenuation is tolerable, are called transmission windows and considered more preferable for 6G purposes [233]. There are massive amounts of spectrum in these windows, ranging from tens to even hundreds of GHz.
@@
    A major challenge is how to handle hardware impairments, such as phase noise and power amplifier non-linearity, as they become more severe at higher frequencies. To a certain extent, a wider SCS can alleviate hardware impairments in multi-carrier-based systems. Another solution is to employ single-carrier waveforms for the operation at very high frequencies since they are more robust against phase noise and other hardware impairments [235]. In terms of transceiver hardware, signal generation and detection are particularly challenging for THz communications [229], and novel solutions are required for cost-, power-, and energy-efficient communications. As THz communication is prone to blockages, a possible solution is to utilize RIS technology, which can reflect the transmitted signals to bypass blockages [236, 237]. With controlled reflections, RISs can also be used to alleviate the distance problem by extending the coverage of THz communications [236, 237]. High power consumption of THz devices can become a limiting factor, highlighting the need for energy-efficient transmission techniques, low-power components, and optimized circuit design. Hybrid beamforming, for instance, offers an excellent balance between power efficiency and performance, making it a potential solution. THz devices also generate a substantial amount of heat, possibly degrading performance and shortening lifespan. This calls for efficient thermal management mechanisms in terms of innovative materials, heat-aware circuit design, and advanced cooling systems.
  • Literature and Future Directions: Mobile THz communication has been one of the major 6G topics since the late 2010s. The THz literature has covered a broad range of topics, such as device technologies, channel modeling, communication designs, networking aspects, application scenarios, regulation/standardization, and experiments [228]. The latest THz research has been comprehensively reviewed in numerous studies [238, 239, 229, 232, 240, 237, 241, 242, 243, 236, 244, 245, 228, 246, 247, 248, 249, 250, 251, 252, 253, 254, 255, 256, 257, 258].
    THz communication systems were reviewed in [238] from the perspective of signal generation, channel modeling, contender technologies, applications, standardization, and future directions. In [239], the THz band was considered for sensing, imaging, and localization applications, especially in the context of 6G use cases. THz ultra-massive MIMO was comprehensively examined in [229] and [244]. Standardization for sub-THz communications in IEEE 802.15.3d was discussed in [232]. 6G THz networks were studied from the full-communication stack perspective in [240], addressing link- and system-level challenges. RIS-assisted THz communication was explored for 6G networks in [237, 236, 252]. In [241], a tutorial was provided on signal processing techniques for THz communications, with an emphasis on ultra-massive MIMO and RIS systems.
@@
IX-A2 OPTICAL WIRELESS COMMUNICATIONS